OpenStudy (anonymous):

A rectangular field is 450 meters long and 300 meters wide. What is the area of the field in square kilometers? Do not round your answer. Be sure to include the correct unit in your answer.

OpenStudy (anonymous):

157.5, 450*300/1000

OpenStudy (anonymous):

157.5 sq km

OpenStudy (anonymous):

It is probably easier if you convert everything into km first.1 Km = 1000m 450m = 450/1000 km = .45 Km 300m = 300/1000 km = .3 km total area = .3 * .45 = .135 sq. km

OpenStudy (anonymous):

you got me i didn't square the 1000 i divied by ((450*300)/(1000^2))=.135sq. km